OpenStudy (anonymous):

Choose the correct simplification of the expression (4x/y)^2

OpenStudy (unklerhaukus):

\[\left(\frac{ab}{c}\right)^n=\frac{a^nb^n}{c^n}\]

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

distribute the square

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

what did you get? (ill confirm if your right)